# JADES: Low Surface Brightness Galaxies at 0.4 < z < 0.8 in GOODS-S

### v1

@@ 1 Introduction
Low surface brightness galaxies (LSBs) are galaxies with a surface brightness dimmer than that of the ground-based night sky, and as a result have been historically difficult to detect and observe. Early research on LSBs suggested a lower limit to a galaxy’s surface brightness (Freeman, 1970). It was then suggested that this was not a physical law, but rather an effect of selection bias due to detection limits at the time (Disney, 1976). Since then, it has been confirmed that galaxies dimmer than the ground-based sky background exist, and constitute a sizeable fraction (up to 60%) of the local galaxy population (McGaugh, 1996; O’Neil & Bothun, 2000; Minchin et al., 2004). We present here a search for LSBs at moderate ($0.4<z<0.8$) redshift using JWST. We describe how these newly observed objects fit into the context of older LSB definitions.
The quantitative definition of a “low surface brightness” galaxy varies from study to study and is primarily based on the detection limits of a given survey. When looking for LSBs, the goal is to find relatively faint and somewhat extended objects. In the past, LSBs have typically been defined as having a $B$-band central surface brightness of $\mu_{0}(B)\geq 23$ mag arcsec-2 (Bothun et al., 1997; Impey & Bothun, 1997). More recent studies have used mean effective surface brightness to consider high surface brightness (HSB) bulges in LSB disks, as relying on central surface brightness can lead to exclusion of such disks (Greco et al., 2018).
Common characteristics of LSBs from past surveys indicate that LSBs are diverse in color, morphology, and environment. LSBs in past surveys have low star formation rates (SFRs) (van den Hoek et al., 2000; Burkholder et al., 2001), are abundant in neutral hydrogen (HI) (Schombert et al., 1992; O’Neil et al., 2004) but deficient in ionized hydrogen (HII) (McGaugh et al., 1995), are metal poor (McGaugh, 1994; de Blok & van der Hulst, 1998b), dust poor (Rahman et al., 2007), and lack molecular gas (CO) (de Blok & van der Hulst, 1998a). Defining LSBs as having a mean effective $g$-band surface brightness of $\bar{\mu}_{\textup{eff}}(g)$ > 24.3 mag arcsec-2, Greco et al. (2018) studied $\sim$700 LSB galaxies in the local Universe with the Hyper Suprime-Cam Subaru Strategic programme (HSC-SSP). They effectively split their sample into blue and red populations. While these LSBs spanned many different environments and morphologies, the blue LSBs in their sample had irregular morphologies with signs of ongoing star formation, while the red LSBs tended to have smooth light profiles well-characterized by a single Sérsic profile, with signs of older stellar populations.
@@
The environments of LSBs are diverse; they are found to be common in clusters (Koda et al., 2015), but they have been observed in all environments (Merritt et al., 2016; Papastergis et al., 2017; Román & Trujillo, 2017; Greco et al., 2018). Compared to the HSB population, LSBs appear to be quenched, and mostly occupy the red sequence (van Dokkum et al., 2015, 2016; Ferré-Mateu et al., 2018; Ruiz-Lara et al., 2018). However, they can be bluer in dense environments (Román & Trujillo, 2017; Zaritsky et al., 2019). This observational evidence, along with predictions from hydrodynamical simulations, suggest that there is not one characteristic evolutionary mechanism alone which is responsible for the formation of LSB galaxies.
@@
Because of the difficulty in finding large samples of LSBs, the majority of studies have been limited to observing those in the local Universe ($z\lesssim 0.15$, Geller et al., 2012; Greene et al., 2022). Prior observational studies have suffered from primarily being ground-based (Rosenbaum & Bomans, 2007; Venhola et al., 2017; Zaritsky et al., 2024), which limits the resolution and available wavelengths, restricting the redshift and luminosity of the observed LSB populations. With the launch of JWST (Gardner et al., 2023), we can now observe LSBs at longer wavelengths with higher resolution and improved sensitivity than before, enabling the calculation of more precise photometric redshifts. Central to this effort is JWST’s Near Infrared Camera (NIRCam) providing a wavelength range of $0.6$ to $5.0\;\mu m$ (Rieke et al., 2023a).
@@ 2 Data and Photometric Catalogue
The HST/ACS mosaics used in this work were produced as part of the Hubble Legacy Fields (HLF) project v2.0 (Illingworth et al., 2016; Whitaker et al., 2019). The JWST/NIRCam data were primarily obtained as part of the JADES (Eisenstein et al., 2023a) programme IDs (PIDs) 1180, 1210, 1286, 1287, 3215, and 6541. In addition, photometric data from the following additional JWST programmes were used: Windhorst IDS GTO programme (PID 1176; Windhorst et al., 2017), Mid-Infrared Instrument (MIRI) Deep Imaging Survey (MIDIS; PID 1283; Östlin et al., 2025), First Reionization Epoch Spectroscopic COmplete Survey (FRESCO; PID 1895; Oesch et al., 2023), JWST Extragalactic Medium-band Survey (JEMS; PID 1963; Williams et al., 2023), Next Generation Deep Extragalactic Exploratory Public (NGDEEP) Survey (PID 2079; Bagley et al., 2024), (PID 2198; Barrufet et al., 2021), PANORAMIC survey (PID 2514; Williams et al., 2025), (ALESS-JWST; PID 2516; Hodge et al., 2021), and BEACON (PID 3990; Morishita et al., 2023).
For a detailed description of the JADES JWST/NIRCam imaging data reduction and mosaicing, consult Tacchella et al. (2023). The JWST/NIRCam source detection is detailed in Robertson et al. (2023). The main steps are stacking six image mosaics (F200W, F277W, F335M, F356W, F410M, and F [^4] W) using the corresponding error images and inverse-variance weighting to produce a single detection image. Within this detection image, a source catalogue is constructed by selecting contiguous regions of greater than 5 pixels (0.15 arcsec) with a signal-to-noise ratio (SNR) of > 3 and applying the SExtractor (Bertin & Arnouts, 1996) deblending algorithm (Bradley et al., 2022). Photometry convolved to the PSF of F [^4] W is then performed at the source centroids in all of the aforementioned photometric bands from HST/ACS and JWST/NIRCam (with the exception of the F430M, F [^4] W, F460M, and F480M bands) with circular ($0.1^{\prime\prime}-0.5^{\prime\prime}$ radius) apertures and elliptical Kron apertures (Kron parameter 1.2, i.e. “small Kron” and Kron parameter 2.5, i.e. “large Kron”). Uncertainties are estimated by placing random apertures across regions of the image mosaics to compute a flux variance. To account for differing depths across the fields, these random apertures are collected into percentiles based on their location’s exposure time, which are then used to determine the sky-noise contribution to the flux uncertainty for each object.
@@ 3 Sample Selection
For details on the derivations of these formulae, consult Graham & Driver (2005).
@@
Motivated by Greco et al. (2018), we choose mean effective surface brightness $\bar{\mu}_{\rm eff}$ (Eqn (4)) as our defining parameter instead of central surface brightness (Eqn (5)) to allow consideration of HSB bulges in extended LSB disks. While the half-light radius $R_{\rm eff}$ can be larger than the object’s physical size, some objects have a fitted $R_{\rm eff}$ that is many times larger than the catalogue semi-major axis for the source, which leads to an underestimation in the mean effective surface brightness. Therefore, we first require that objects have a fitted effective radius less than 1.5 times their catalogue semi-major axis ($R_{\rm eff}<1.5a$) to avoid this underestimation. Since we want to focus on somewhat extended LSBs, we then require that our objects need an effective radius of at least 0.18 arcsec.
@@
Finally, we visually inspect each of these objects to identify false positives and ensure our sample is comprised of actual faint and extended galaxies. These false positives can include tidal tails, shredded galaxy subcomponents, and bits of diffraction spikes from bright stars. False-colour RGB images from the F115W, F200W, and F277W filters of many of the largest LSBs (by fitted $R_{\rm eff}$) in this sample are shown in Figure 4 at fixed 50 x 50 pixel (1.5 x 1.5 arcsec) panels. After these visual cuts, we are left with 57 galaxies in our final LSB sample, with photometric redshifts in the range of $0.4<z<0.8$. Our analysis suggests JADES data cannot detect LSBs ($\bar{\mu}_{\rm eff}(\rm F200W)>24\;mag\;arcsec^{-2}$) past $z\sim 0.8$ in the way we have defined them. Results from pyimfit for these 57 LSBs are reported in Table 1.
Figure 4: False-colour RGB (F277W/F200W/F115W) images for the 20 largest LSBs in the sample, by effective radii. Each cutout is 50 x 50 pixels (1.5 x 1.5 arcsec), and each object is normalised by its individual cutout’s brightest pixel. Each object’s JADES catalogue ID and photometric redshift are shown in the lower left corners, while the physical scale corresponding to 0.3 arcsec is provided in the top right corners.
@@ 4.2 Spatial Distribution
The LSBs are not uniformly distributed; they generally are more abundant in the deeper parts of the field, which is a selection effect. Most of these LSBs appear in regions of GOODS-S with deep coverage: out of the 57 galaxies in our sample, 18 can be found in the JADES Deep field, 11 in the NGDEEP survey, 10 in the JADES Origins Field (JOF; Eisenstein et al., 2023b), and 10 in JADES PID 1287, all of which contain the deepest data in this field. This implies that survey depth is a strong driver of finding LSBs at higher redshift.
@@ 4.3 Number Density
For our LSB sample, we can do a rough calculation of the number density of LSBs in GOODS-S. The total survey area of the JADES GOODS-S field considered here is $\sim$200 arcmin2. Using the comoving distance between $z\sim 0.4$ and $z\sim 0.8$ to approximate the total volume searched, we estimate the comoving volume of the area searched to be $\sim 43,910$ Mpc3. Using this comoving volume estimation, we find that for our 57 LSBs, the number density of these objects in the way we’ve defined them as $\mu_{\rm eff}(\rm F200W)<24$ mag arcsec-2 is on the order of $\sim 10^{-3}$ Mpc-3. This is consistent with the LSB density values reported in Greene et al. (2022) for the range of $M_{*}\sim 10^{7}-10^{8}M_{\odot}$, which is roughly the range of stellar masses for our LSB sample (see Figure 9).
@@ 4.4 Surface Brightness Profiles (Multi-Band)
As described in Section 3, the LSBs in our sample underwent an initial round of surface brightness fitting with pyimfit in the F200W band. This first round of fitting is run on all the objects in our catalogue with SNR >10>10 in F200W to measure relative surface brightness, and PSF broadening is not accounted for. The initial fits are for picking out the faintest, most extended objects past $z_{\rm phot}>0.4$. We further perform a more deliberate, PSF-convolved, accurate Sérsic profile fit in multiple filters, for which we employ pysersic (Pasha & Miller, 2023). pysersic is more recently developed in contrast to pyimfit, utilizes Bayesian inference for speed, and is specialised for single-component Sérsic fits. The purpose of this second round of fitting is to gather accurate Sérsic profile parameters in different filters for each of our LSBs, as well as measuring the surface brightness in multiple filters that take PSF convolution into account. At a base level, pysersic operates much in the same way as pyimfit does, with similar fit results as what is shown in Figure 2 for pyimfit. A direct comparison figure is provided in Appendix A.
@@ 5.2 Stellar Mass and Dust Attenuation
The first photometrically-derived quantities we want to study from bagpipes are stellar mass $M_{*}$ and dust attenuation $A_{V}$. Stellar mass can highlight the differences in star formation between the LSBs and HSBs, while dust attenuation can tell us if light is being obscured, and if dust could be another explanation for why LSBs are low surface brightness. Histogrammes of stellar mass (left) and V-band dust attenuation $A_{V}$ (right) are shown in Figure 9, with colouring of samples consistent with Figures 11 and 12.
In general, we can see that the left panel of Figure 9 validates our choice of F200W magnitude as a prior for stellar mass, as all three samples lie in the dwarf ($M_{*}<10^{9}M_{\odot}$) regime. While there is significant overlap between the distributions of stellar mass for the three samples, Figure 9 shows a trend of increasing stellar mass from the $z<0.4$ LSBs to the $z>0.4$ LSBs to the HSBs, being partly a selection effect of the brightness criterion, with most of the $z>0.4$ LSBs have higher stellar masses than the $z<0.4$ LSBs.
@@
Figure 10: Size-mass relation for all three of our selected galaxy samples. The best fit effective radius from pyimfit is plotted as $\log(R_{\rm eff}/\rm kpc)$ against the posterior median stellar mass from bagpipes is plotted as $\log(M_{*}/M_{\odot})$. The best fit size-mass relation for early type (dotted line) and late (solid line) type galaxies from van der Wel et al. (2014) are shown as black lines for their $0.5<z<1.0$ redshift bin. Sample colouring is consistent with Figure 9.
@@
With a measure of size from the best-fit effective radius from pyimfit (see Section 3) and an estimate of stellar mass from bagpipes, we can compare our galaxy samples with previously fitted size-mass relations to see how they all fit into the general population. We use these two quantities to show a size-mass relation in Figure 10, and overplot the best-fit lines for galaxies $0.5<z<1$ from the 3D-HST+CANDELS survey reported in van der Wel et al. (2014) for both early and late type galaxies (fits shown in their Figure 5 and values reported in their Table 1). The late-type galaxy fit best describes the size-mass relation for all three of our samples.
From Figure 10, it is shown that most of the $0.4<z<0.8$ LSBs sample lies above the van der Wel et al. (2014) fits at larger effective radii, while most of the HSB sample lies below the fits with smaller radii at similar stellar mass. This is an effect of our selection process, as the HSBs were selected based off of having similar total magnitude but higher surface brightness, which results in the HSB sample being more compact than the LSB sample. The $z<0.4$ LSB sample lies on either side of the van der Wel et al. (2014) fits, but at lower stellar masses on average. No particular sample is shown to be major outliers in the size-mass relation for the general galaxy population.

### v2

@@ 1 Introduction
Low surface brightness galaxies (LSBs) are galaxies with a surface brightness dimmer than that of the ground-based night sky, and as a result have been historically difficult to detect and observe. Early research on LSBs suggested a lower limit to a galaxy’s surface brightness (Freeman, 1970). It was then suggested that this was not a physical law, but rather an effect of selection bias due to detection limits at the time (Disney, 1976). Since then, it has been confirmed that galaxies dimmer than the ground-based sky background exist, and constitute a sizeable fraction (up to 60%) of the local galaxy population (McGaugh, 1996; O’Neil and Bothun, 2000; Minchin et al., 2004). We present here a search for LSBs at moderate ($0.4<z<0.8$) redshift using JWST. We describe how these newly observed objects fit into the context of older LSB definitions.
The quantitative definition of a “low surface brightness” galaxy varies from study to study and is primarily based on the detection limits of a given survey. When looking for LSBs, the goal is to find relatively faint and somewhat extended objects. In the past, LSBs have typically been defined as having a $B$-band central surface brightness of $\mu_{0}(B)\geq 23$ mag arcsec-2 (Bothun et al., 1997; Impey and Bothun, 1997). More recent studies have used mean effective surface brightness to consider high surface brightness (HSB) bulges in LSB disks, as relying on central surface brightness can lead to exclusion of such disks (Greco et al., 2018).
Common characteristics of LSBs from past surveys indicate that LSBs are diverse in color, morphology, and environment. LSBs in past surveys have low star formation rates (SFRs) (van den Hoek et al., 2000; Burkholder et al., 2001), are abundant in neutral hydrogen (HI) (Schombert et al., 1992; O’Neil et al., 2004) but deficient in ionized hydrogen (HII) (McGaugh et al., 1995), are metal poor (McGaugh, 1994; de Blok and van der Hulst, 1998b), dust poor (Rahman et al., 2007), and lack molecular gas (CO) (de Blok and van der Hulst, 1998a). Defining LSBs as having a mean effective $g$-band surface brightness of $\bar{\mu}_{\textup{eff}}(g)$ > 24.3 mag arcsec-2, Greco et al. (2018) studied $\sim$700 LSB galaxies in the local Universe with the Hyper Suprime-Cam Subaru Strategic programme (HSC-SSP). They effectively split their sample into blue and red populations. While these LSBs spanned many different environments and morphologies, the blue LSBs in their sample had irregular morphologies with signs of ongoing star formation, while the red LSBs tended to have smooth light profiles well-characterized by a single Sérsic profile, with signs of older stellar populations.
@@
The environments of LSBs are diverse; they are found to be common in clusters (Koda et al., 2015), but they have been observed in all environments (Merritt et al., 2016; Papastergis et al., 2017; Román and Trujillo, 2017; Greco et al., 2018). Compared to the HSB population, LSBs appear to be quenched, and mostly occupy the red sequence (van Dokkum et al., 2015, 2016; Ferré-Mateu et al., 2018; Ruiz-Lara et al., 2018). This observational evidence, along with predictions from hydrodynamical simulations, suggest that there is not one characteristic evolutionary mechanism alone which is responsible for the formation of LSB galaxies.
@@
Because of the difficulty in finding large samples of LSBs, the majority of studies have been limited to observing those in the local Universe ($z\lesssim 0.15$, Geller et al., 2012; Greene et al., 2022). Prior observational studies have suffered from primarily being ground-based (Rosenbaum and Bomans, 2007; Venhola et al., 2017; Zaritsky et al., 2024), which limits the resolution and available wavelengths, restricting the redshift and luminosity of the observed LSB populations. With the launch of JWST (Gardner et al., 2023), we can now observe LSBs at longer wavelengths with higher resolution and improved sensitivity than before, enabling the calculation of more precise photometric redshifts. Central to this effort is JWST’s Near Infrared Camera (NIRCam) providing a wavelength range of $0.6$ to $5.0\;\mu m$ (Rieke et al., 2023a).
@@ 2 Data and Photometric Catalogue
The HST/ACS mosaics used in this work were produced as part of the Hubble Legacy Fields (HLF) project v2.0 (Illingworth et al., 2016; Whitaker et al., 2019). The JWST/NIRCam data were primarily obtained as part of the JADES (Eisenstein et al., 2023b) programme IDs (PIDs) 1180, 1210, 1286, 1287, 3215, and 6541. In addition, photometric data from the following additional JWST programmes were used: Windhorst IDS GTO programme (PID 1176; Windhorst et al., 2017), Mid-Infrared Instrument (MIRI) Deep Imaging Survey (MIDIS; PID 1283; Östlin et al., 2025), First Reionization Epoch Spectroscopic COmplete Survey (FRESCO; PID 1895; Oesch et al., 2023), JWST Extragalactic Medium-band Survey (JEMS; PID 1963; Williams et al., 2023), Next Generation Deep Extragalactic Exploratory Public (NGDEEP) Survey (PID 2079; Bagley et al., 2024), (PID 2198; Barrufet et al., 2021), PANORAMIC survey (PID 2514; Williams et al., 2025), (ALESS-JWST; PID 2516; Hodge et al., 2021), and BEACON (PID 3990; Morishita et al., 2023).
For a detailed description of the JADES JWST/NIRCam imaging data reduction and mosaicing, consult Tacchella et al. (2023). The JWST/NIRCam source detection is detailed in Robertson et al. (2023). The main steps are stacking six image mosaics (F200W, F277W, F335M, F356W, F410M, and F [^4] W) using the corresponding error images and inverse-variance weighting to produce a single detection image. Within this detection image, a source catalogue is constructed by selecting contiguous regions of greater than 5 pixels (0.15 arcsec) with a signal-to-noise ratio (SNR) of > 3 and applying the SExtractor (Bertin and Arnouts, 1996) deblending algorithm (Bradley et al., 2022). Photometry convolved to the PSF of F [^4] W is then performed at the source centroids in all of the aforementioned photometric bands from HST/ACS and JWST/NIRCam (with the exception of the F430M, F [^4] W, F460M, and F480M bands) with circular ($0.1^{\prime\prime}-0.5^{\prime\prime}$ radius) apertures and elliptical Kron apertures (Kron parameter 1.2, i.e. “small Kron” and Kron parameter 2.5, i.e. “large Kron”). Uncertainties are estimated by placing random apertures across regions of the image mosaics to compute a flux variance. To account for differing depths across the fields, these random apertures are collected into percentiles based on their location’s exposure time, which are then used to determine the sky-noise contribution to the flux uncertainty for each object.
@@ 3 Sample Selection
For details on the derivations of these formulae, consult Graham and Driver (2005).
@@
Motivated by Greco et al. (2018), we choose mean effective surface brightness $\bar{\mu}_{\rm eff}$ (Eqn (4)) as our defining parameter instead of central surface brightness (Eqn (5)) to allow consideration of HSB bulges in extended LSB disks. While the half-light radius $R_{\rm eff}$ can be larger than the object’s physical size, some objects have a fitted $R_{\rm eff}$ that is many times larger than the catalogue semi-major axis for the source, which leads to an underestimation in the mean effective surface brightness. Therefore, we first require that objects have a fitted effective radius less than 1.5 times their catalogue semi-major axis ($R_{\rm eff}<1.5a$) to avoid this overestimation. Since we want to focus on somewhat extended LSBs, we then require that our objects need an effective radius of at least 0.18 arcsec.
@@
Finally, we visually inspect each of these objects to identify false positives and ensure our sample is comprised of actual faint and extended galaxies. These false positives can include tidal tails, shredded galaxy subcomponents, and bits of diffraction spikes from bright stars. False-colour RGB images from the F115W, F200W, and F277W filters of many of the largest LSBs (by fitted $R_{\rm eff}$) in this sample are shown in Figure 4 at fixed 50 $\times$ 50 pixel (1.5 $\times$ 1.5 arcsec) panels. After these visual cuts, we are left with 57 galaxies in our final LSB sample, with photometric redshifts in the range of $0.4<z<0.8$. Our analysis suggests JADES data cannot detect LSBs ($\bar{\mu}_{\rm eff}(\rm F200W)>24\;mag\;arcsec^{-2}$) past $z\sim 0.8$ in the way we have defined them. Results from pyimfit for these 57 LSBs are reported in Table 1.
Figure 4: False-colour RGB (F277W/F200W/F115W) images for the 20 largest LSBs in the sample, by effective radii. Each cutout is 50 $\times$ 50 pixels (1.5 $\times$ 1.5 arcsec), and each object is normalised by its individual cutout’s brightest pixel. Each object’s JADES catalogue ID and photometric redshift are shown in the lower left corners, while the physical scale corresponding to 0.3 arcsec is provided in the top right corners.
@@ 4.2 Spatial Distribution
The LSBs are not uniformly distributed; they generally are more abundant in the deeper parts of the field, which is a selection effect. Most of these LSBs appear in regions of GOODS-S with deep coverage: out of the 57 galaxies in our sample, 18 can be found in the JADES Deep field, 11 in the NGDEEP survey, 10 in the JADES Origins Field (JOF; Eisenstein et al., 2023a), and 10 in JADES PID 1287, all of which contain the deepest data in this field. This implies that survey depth is a strong driver of finding LSBs at higher redshift.
@@ 4.3 Number Density
For our LSB sample, we can do a rough calculation of the number density of LSBs in GOODS-S. The total survey area of the JADES GOODS-S field considered here is $\sim$200 arcmin2. Using the comoving distance between $z\sim 0.4$ and $z\sim 0.8$ to approximate the total volume searched, we estimate the comoving volume of the area searched to be $\sim 43,910$ Mpc3. Using this comoving volume estimation, we find that for our 57 LSBs, the number density of these objects in the way we’ve defined them as $\mu_{\rm eff}(\rm F200W)>24$ mag arcsec-2 is on the order of $\sim 10^{-3}$ Mpc-3. This is consistent with the LSB density values reported in Greene et al. (2022) for the range of $M_{*}\sim 10^{7}-10^{8}M_{\odot}$, which is roughly the range of stellar masses for our LSB sample (see Figure 9).
@@ 4.4 Surface Brightness Profiles (Multi-Band)
As described in Section 3, the LSBs in our sample underwent an initial round of surface brightness fitting with pyimfit in the F200W band. This first round of fitting is run on all the objects in our catalogue with SNR >10>10 in F200W to measure relative surface brightness, and PSF broadening is not accounted for. The initial fits are for picking out the faintest, most extended objects past $z_{\rm phot}>0.4$. We further perform a more deliberate, PSF-convolved, accurate Sérsic profile fit in multiple filters, for which we employ pysersic (Pasha and Miller, 2023). pysersic is more recently developed in contrast to pyimfit, utilizes Bayesian inference for speed, and is specialised for single-component Sérsic fits. The purpose of this second round of fitting is to gather accurate Sérsic profile parameters in different filters for each of our LSBs, as well as measuring the surface brightness in multiple filters that take PSF convolution into account. At a base level, pysersic operates much in the same way as pyimfit does, with similar fit results as what is shown in Figure 2 for pyimfit. A direct comparison figure is provided in Appendix A.
@@ 5.2 Stellar Mass and Dust Attenuation
The first photometrically-derived quantities we want to study from bagpipes are stellar mass $M_{*}$ and dust attenuation $A_{V}$. Stellar mass can highlight the differences in star formation between the LSBs and HSBs, while dust attenuation can tell us if light is being obscured, and if dust could be another explanation for why LSBs are low surface brightness. Histogrammes of stellar mass (left) and V-band dust attenuation $A_{V}$ (right) are shown in Figure 9.
In general, we can see that the left panel of Figure 9 validates our choice of F200W magnitude as a prior for stellar mass, as all three samples lie in the dwarf ($M_{*}\lesssim 10^{8}M_{\odot}$) regime. While there is significant overlap between the distributions of stellar mass for the three samples, Figure 9 shows a trend of increasing stellar mass from the $z<0.4$ LSBs to the $z>0.4$ LSBs to the HSBs, being partly a selection effect of the brightness criterion, with most of the $z>0.4$ LSBs have higher stellar masses than the $z<0.4$ LSBs.
@@
[…]
@@
[…]
From Figure 10, it is shown that most of the $0.4<z<0.8$ HSB sample lies along the ELVES fit within uncertainty, while most of the $0.4<z<0.8$ LSB sample lies above the fit with larger radii at similar stellar mass. The LSBs being more extended than the HSBs at fixed stellar mass is an effect of our selection process, as the HSBs were selected based off of having similar total magnitude but higher surface brightness, which results in the HSB sample being more compact than the LSB sample. What Figure 10 reveals is that the higher redshift LSBs are outliers in the size-mass relation for local dwarf galaxies, whereas the majority of the HSB sample are not outliers. This supports the idea of LSBs having higher spin parameters early in their evolution (Pérez-Montaño et al., 2022), leading to extended morphologies that result in low surface brightness. Roughly half of the $z<0.4$ LSB sample lies within the uncertainty of the ELVES fit, while the other half are extended outliers like the $0.4<z<0.8$ LSBs.